If the sides of a square are lengthened by 8cm, the area becomes 169cm^2 find the length of a side of the orginal square. This one has me stumped and am not sure how to solve it.
Answer by mananth(16946)

let the side be x
new side = x+8
(x+8)(x+8)=225
(x+8)^2=15^2
Take square root
x+8= 15
x= 15-8
x=7 cm